OpenStudy (anonymous):

express y as a function of x. what is the domain? a) log 3 + log y = log (x+2) - log x this is what i did: log 3 + log y = log (x+2) - log x log y = log (x+2) - log x - log 3 log y = log (x+2) - log (3x) log y = log [(x+2)/(3x)] y = (x+2)/(3x) .. is this right? or did i do something wrong.. my friend did it another way.. so idk

OpenStudy (lgbasallote):

yes you're right

OpenStudy (anonymous):

just to clarify.. if i have log on both the right and left side... do i just "remove it" ??

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

however if you have something like \[\log y = x\] it will become \[y = 10^x\]

OpenStudy (anonymous):

okay. so i'm given another equation where log 4y = x + log 4... would this become log 4y = 10 x + log 4 ? then log 4y = log 40x ... ? :S

OpenStudy (lgbasallote):

\[\huge x \implies x(1) \implies x(\log 10)\]

OpenStudy (anonymous):

SO... it would be log 4y = x(log10) + log 4

OpenStudy (anonymous):

?*

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

you can combine it further

OpenStudy (anonymous):

log 4y = x(log40) ?

OpenStudy (lgbasallote):

no..

OpenStudy (lgbasallote):

\[x \log 10 \implies \log 10^x\]

OpenStudy (anonymous):

uhm.. \[\log 4y = \log 10^{x} + \log 4\]\[\log 4y = \log (10^{x} + 4)\]

OpenStudy (lgbasallote):

10^x + 4?

OpenStudy (anonymous):

40^x ?

OpenStudy (anonymous):

my bad.. i meant 40^x

OpenStudy (lgbasallote):

still no

OpenStudy (lgbasallote):

\[\huge a^m \times b \ne (ab)^m\]

OpenStudy (lgbasallote):

10^x * 4 is 10^x *4....you cant do anything about it

OpenStudy (anonymous):

... oh my gosh.. sorry! okay so it would then be log 4y = log(10^x(4)) 4y = (10^x(4)) y = 10^x

OpenStudy (lgbasallote):

yes!!

OpenStudy (anonymous):

i am so sorry... ! im just making this a lot harder than it should be... thank you so much !